Belgium is a small country which produces and consumes
gooseberries. The quantity demand of gooseberries is Qd= 120-4P, and
quantity supply is Qs= 2P in Belgium. If world price of gooseberries is
Rs 15 then Belgium will export or import gooseberries and by how
much?
1
Expert's answer
2021-07-20T09:58:48-0400

Given

 Qd= 120-4P

 Qs= 2P

P=Rs 15


"Qd= 120-4P\\ Qd= 120-4(15)=60"


"Qs= 2P\\ Qs= 2(15)=30"


since the quantity demanded is more than the quantity supplied. Belgium will import gooseberries to meet the deficit.


The quantity imported will be:

"60-30=30"

Therefore to meet the demand, Belgium is required to import 30 goose berries.